Find the center and radius of x^2+y^2-6x-2y-6=0
This is non from a text book

Answer by fractalier(6550)

From x^2+y^2-6x-2y-6=0 we complete the squares...
x^2 - 6x + ___ + y^2 - 2y + ___ -6 = 0 + ___ + ___
x^2 - 6x + 9 + y^2 - 2y + 1 = 6 + 9 + 1
(x - 3)^2 + (y - 1)^2 = 16
center at (3, 1) and radius = 4